HRMS, 28th Colloquium, Dijon, France, Aug. 28 - Sept. 1, 2023

TEMPERATURE DEPENDENCE OF THE ABSORPTION OF THE R(6) MANIFOLD OF THE 2NU3 BAND OF METHANE IN AIR IN SUPPORT OF

THE MERLIN MISSION

S. VASILCHENKO, S. KASSI, A. CAMPARGUE, \underline{D. MONDELAIN} , Univ. Grenoble Alpes, CNRS, LIPhy, 38000 Grenoble, France; \underline{T. DELAHAYE} , R. ARMANTE, H. TRAN, Laboratoire de Météorologie Dynamique, IPSL, CNRS,

Ecole polytechnique, Université Paris-Saclay, Palaiseau, France

Fifty nine high sensitivity spectra of the R(6) manifold of the 2ν_{3} band of methane in air, near 1.64 μm, have been recorded in support of the MERLIN mission. For this purpose, a cavity ring down spectrometer (CRDS) with a spectrally narrowed and stable (sub-kHz) laser source was coupled to a temperature regulated high-finesse optical cavity. The frequency scale of each spectrum was accurately determined from measurements of the beat note between a part of the laser light and the closest tooth of a frequency comb referenced to a rubidium clock. Series of spectra were recorded between 243 and 313 K with a 10 K temperature step. For each series, total pressure values of 50, 100, 250, 500 and 750 Torr were adopted. A multi-spectrum fitting procedure with the Hartmann-Tran (HT) line profile, including the first-order linemixing parameter, has been used to derive the spectroscopic parameters for each of the six R(6) components, along with the temperature dependence of the line-shape parameters. The results show that the fitted model is able to reproduce the experimental spectra with a relative precision better than 0.2% for the entire R(6) manifold spectral region and better than 0.05% at the ON-line position of the MERLIN mission for the 250, 500 and 750 Torr spectra. The relative precision increases to 0.3% and the residuals at the ON-line position to 0.1% when including the 50 and 100 Torr spectra. Comparisons with ground-based atmospheric measurements show that these data significantly improve the modeling of methane absorption in this spectral region. We propose here a new complete line list of the methane spectrum in the region of the R(6) manifold allowing reducing notably the residuals at the ON-line position of the MERLIN mission.

p-number: p042  Submitted on Tue, 04 Jul 2023 12:04:53 +0200